Calculate the momentum. of a tody which is moving with a velocity of 72km/es and having 8 mass 2500 g


1
Expert's answer
2021-07-05T08:43:11-0400
"p=mv\\p=\frac{72}{3.6}(2.5)=50\frac{kg\cdot m}{s}"